3. (5) Solve the initial-value problem: ( 20) 106 x, subject to x(0) =
Answer #1

Sol. x=(8%) * ; *(0) = ). The first thing that we need to calculate is the eigenvalues for the matrix A = (6 %): det (A-NI)Me - Gule 1e 0 B Eigenvectos corresponding to the eigenvalue G is : cl Bil 7 o 6 ( B2 / 1 + 1) - D = 2P) = 6B=> 48, = 0 37 P,an method : 7=66)* ; 260)=(%) (*):( 32) () (7) 2x del = 2x, - 0 dX de = 6X2 from 6; we get dx = adt 24 Integrate on both side- len az = 67 + Ao where Ao is ambitioning O constant ng = est+to = eho eft = Geft where G=eto [x2= Geft ] put X (0) = 6 ; we